OpenStudy (anonymous):

Convert log110 to an expression containing only natural logarithms

OpenStudy (anonymous):

change of base formula says \[log_b(x)=\frac{log_a(x)}{log_a(b)}\] but i cannot read exactly what you wrote. your answer will look like \[\frac{ln(x)}{ln(10)}\] if your base is ten.

OpenStudy (anonymous):

Thanks, the question is really Log10(110) (10 being th base) So what is the x value on the top of the fraction [ln(x)]?

OpenStudy (anonymous):

ok change of base formula: on top goes log of input, on bottom goes log of base.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the base here is ten, so you use \[log_10(110)=\frac{ln(110)}{ln(10)}\]

OpenStudy (anonymous):

this is why fancy calculators only have the natural log. you can always convert. it is also how you solve \[b^x=A\] for x

OpenStudy (anonymous):

Thanks alot my good man.

OpenStudy (anonymous):

welcome